17 The table below shows the daily costs of Cornys Corn Cobs. Cornys sells corn by the dozen in a perfectly competitive marc. In the short run, assuming nothing else changes, Cornys should O shut down, because the market price is above the AVC. O
Answer #1

17. (a) In the short run, Corny's produce should be 70 dozen ears of corns per day.

The production should take place where MC=P, which is at where Q=70 and Q=80. Since at both production levels, we have ATC>P, the minimization of loss would be where Q=70.

(b) Profit of loss per day would be $-7.

Since price is less than the AC, there will be loss, not profit. Also, the amount would be T= (P - ATC) or T= 2.2-2.3)* 70 or \pi = -7 .

(c) The correct option would be

  • produce the same quantity of corn per day.

Yet firm is making loss, the firm should produce in the short run if the price is above the AVC at the production level. This would mean that the firm is still covering its variable costs, and should continue to produce in the short run. In the long run however, if these circumstances pertains, then the firm should leave the market.

(d) The correct option would be

  • shut down, because the market price is below the AVC.

In the short run, the firm should shut down if it can not cover its variable costs, which would be the case is price goes below the AVC. As can be seen, the price of $1.30 is below all the AVCs, and hence the firm should shut down.